A cheetah can accelerate from rest to a speed of 30.0 m/s in 7.00 s. What is its acceleration in m/s2? Use Two Significant Figures in your answer


1
Expert's answer
2021-01-24T16:29:46-0500
"a=\frac{v}{t}\\a=\frac{30}{7}=4.29\frac{m}{s}"